# LASeR: Learning to Adaptively Select Reward Models with Multi-Armed Bandits

### v2

@@ 6 Conclusion
Limitations and Broader Impacts
Limitations.
While LASeR demonstrates strong empirical performance and generalization across tasks, policy models, and RMs, it still has some limitations. First, like all work using RMs to improve LLMs, it relies on the availability of a high-quality pool of RMs, and downstream performance may degrade if the pool contains only weak or biased models. Additionally, although the bandit-based approach mitigates reliance on a single RM, it does not explicitly address issues of safety, fairness, or bias in RM selection and consider potential trade-offs between safety and task effectiveness when choosing an RM. Finally, besides the MAB formulation, other methods such as reinforcement learning-based RM selection might be applicable. This could allow for more flexible RM selection policies, but would introduce substantial optimization and stability challenges, especially when dealing with non-stationary and noisy RM feedback. We consider these directions as promising for future research.
Broader Impact.
LLMs have been shown to reflect stereotypes, biases, and other negative traits contained in their pretraining data (Weidinger et al., 2021). Consequently, finetuned LLMs (including those trained with LASeR) may also exhibit such undesirable traits in their generations during inference or training and exhibit the same potential for misuse as any other finetuned model. While prior work has made some headway in detecting such harmful content generated by LLMs (Inan et al., 2023), considerable research effort is needed in mitigating bias in LLMs. Conceptually, classifiers that detect risky, harmful, or biased content in the text can also be used as an additional RM in LASeR’s training to reinforce avoiding bias via preference optimization. However, we do not study this in our work and leave it to future work to explore these directions.
@@ Acknowledgements
We thank Justin Chih-Yao Chen for his feedback on the figures. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, the Accelerating Foundation Models Research Program, and NSF CAREER Award 1846185. The views contained in this article are those of the authors and not of the funding agency.
@@ A.1 Experimental Setting
Training Setup.
We provide the training details for our method and baselines as follows:
@@
  • Batch size: We fine-tune the model using a learning rate of $65e\!-\!$ 6 and a batch size of 16. Noted that we experimented with different batch sizes to evaluate their impact. Using a batch size of 1 yielded comparable performance to a batch size of 16 but was significantly less efficient in training the LLM due to the increased computational overhead. Based on this, we opted to use a batch size of 16 for a better trade-off between performance and efficiency. For datasets like Wildchat, which contain clearly defined and diverse categories, we structure the batches such that each batch consists of data belonging to a single category. This setup minimizes the risk of mismatches between the RM and the batch data, as each RM is evaluated on its most relevant data category. For reasoning datasets where such predefined categories do not exist, shuffling the data during training ensures diverse data and a good training signal for LLM within each batch.
@@
To test the hypothesis that the log probabilities of the chosen and rejected responses from the policy model are generally close in value, with the chosen response tending to have a slightly higher log probability, we conduct an experiment using Llama-3-8B and RewardBench dataset. For each instance, we calculated the reward difference by subtracting the log probability (reward) of the chosen response from that of the rejected response. Formally, given a policy model $\pi_{0}$, a prompt $x_{i}$ and a pair of chosen response $y_{i}^{w}$ and rejected response $y_{i}^{l}$ from RewardBench data, we calculate the reward difference as $\log\pi_{0}(y_{i}^{w}\mid x_{i})-\log\pi_{0}(y_{i}^{l}\mid x_{i})$. As shown in Fig. 5, the distribution is centered near zero, indicating that the values are often close. However, the mass is skewed slightly to the right, suggesting that the chosen response tends to have a higher log probability than the rejected one in the majority of cases. This suggests the policy model generally has a reasonable ranking, but benefits from DPO’s maximization of the margin.
@@
Datasets. For StrategyQA, GSM8K, and MMLU, we divided each dataset into training and test sets. The model is fine-tuned on the training set and dev set and then evaluated on the test set. For WildChat, the dataset was split into a 70/10/20 ratio for training, development, and testing. Following Zhao et al. [2024], prompt categorization is done using a small off-the-shelf classifier. [^5] Link: https://huggingface.co/valpy/prompt-classification For LongBench, we subsample 5K examples for three tasks: multi-document QA, summarization, and few-shot learning. Each category was split into a 70/10/20 ratio, and the bandit model was trained and validated on the training and development sets and then tested on the test set. We report the detailed number of instances for train, development, and test sets in Table 3.
@@
Additional Results and Analysis. In Appendix B, we present additional analyses on the statistical stability of LASeR and baselines across multiple runs (Table 4). We show that the RM selected by LASeR adapts to input queries (Fig. 6) and aligns well with RewardBench ground truth rankings (Fig. 8). We also show that MAB rewards converge over training (Fig. 9), and that the learned MAB policy can be reused to improve cold-start training (Table 5). Detailed comparisons against individual RMs used in isolation (Tables 6 and 7) further confirm that LASeR consistently outperforms all single-RM baselines. In addition, we show that LASeR maintains strong performance across different sets of RMs (Table 8).
In Appendix C, we evaluate the generalization of LASeR across multiple dimensions: different bandit algorithms (Table 9), base models and RMs (Tables 10 and 12), number of RMs (Fig. 11), training loss functions (Table 13), and out-of-distribution tasks (Table 14). We also show that LASeR can successfully select evaluation metrics tailored to domain-specific reasoning tasks [Golovneva et al., 2022, Prasad et al., 2023] (Table 15). Finally, LASeR demonstrates robustness to both noisy RMs and underperforming reward signals (Figs. 12 and 13).
@@ A.2 Details of Bandit Algorithms
Algorithm for Sec. 3.3. We provide the detailed algorithm for Sec. 3.3 in Algorithm 1.
@@
1:  Input: LLM $\mathcal{M}$, reward models =$\mathcal{R}=\{R_{1},R_{2},\dots,R_{K}\}$, dataset $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$, bandit algorithm (LinUCB)
@@
The score for arm $a_{t}$ is updated based on the observed normalized reward ^$-\hat{\mathcal{L}}^{m}(t)$ and the probability $p_{k}(t)$ of selecting that arm:
@@
MAB reward normalization. To maintain a consistent scale and magnitude of MAB rewards across training, we apply scaled rewards based on the quantiles of the reward history, following the method outlined by Graves et al. [2017]. Let $L=\{-\mathcal{L}^{m}(1),\dots,-\mathcal{L}^{m}(t-1)\}$ represent the unscaled reward history up to time step $t$. This history’s lower and upper quantiles are denoted as $q_{t}^{lo}$ and $q_{t}^{hi}$, respectively. We set $q_{t}^{\text{lo}}$ and $q_{t}^{\text{hi}}$ to be $20^{th}$ and $80^{th}$ quantiles. The scaled reward, ^$-\hat{\mathcal{L}}^{m}(t)$, becomes:
^$-\hat{\mathcal{L}}^{m}(t)=\begin{cases}0&\text{if }-\mathcal{L}^{m}(t)<q_{t}^{\text{lo}}\\ 1&\text{if }-\mathcal{L}^{m}(t)>q_{t}^{\text{hi}}\\ \frac{-{\mathcal{L}}^{m}(t)-q_{t}^{\text{lo}}}{q_{t}^{\text{hi}}-q_{t}^{\text{lo}}}&\text{otherwise}.\end{cases}$
@@ Appendix B Additional Empirical Results
Analysis on Statistical Significance of LASeR.
While our main results are based on single runs, we note that this is standard practice in iterative preference optimization setups, including recent works such as Pang et al. [2024], where large-scale LLM training limits the feasibility of multi-seed replication across all benchmarks. Nonetheless, to provide some insights regarding the statistical significance of LASeR, we report the mean and standard deviation over 5 seeds on reasoning tasks. We compare LASeR against the Best RM and RM Agreement Ensemble (identical random seeds to ensure fair comparison), which are the most competitive across our benchmarks. Table 4 shows that LASeR consistently outperforms the baselines across runs, with standard deviations below 0.3% (lower than RM Agreement Ensemble), showing stable performance. We believe these consistent gains across tasks and models support the claim that LASeR is not only superior but also reliable in expectation.
@@
LASeR’s Selected RM Adjusts to the Query.
Fig. 6 shows the relative utilization rates of each arm (i.e., RM) of the bandit on WildChat. We observed vastly different RM utilization rates depending on the underlying query within the same dataset. On queries requiring creativity in LLM responses, we find that Olmo and Eurus RMs are utilized about $20\%$ more often than Qwen RM, despite Qwen RM being ranked higher on RewardBench. This can be explained by the fact that the Qwen RM largely underperforms on the “chat” subsplit of RewardBench (behind Olmo and Eurus by nearly 40 points in chat score). On the other hand, Qwen RM is used roughly half the time for user prompts involving math, while Olmo and Eurus are used sparingly. This is consistent with Qwen RM’s ranking on the “reasoning” split of RewardBench, outperforming Eurus and Olmo RMs by 15-20 points. Note that LASeR automatically deduces these relative rankings of RMs and uses them depending on the underlying query without having access to the RewardBench leaderboard. Therefore, RM utilization of LASeR can serve as an analysis tool for future work when assessing performance on untested domains.
@@
To further validate our findings, we conducted an additional experiment on RewardBench data. In this setup, we used the MAB trained on Wildchat to select the most appropriate RM for each query. This selected RM was then used to choose the preferred response, which we compared against the ground-truth preferences in RewardBench. We also report the accuracy of each RM in selecting the correct preferred responses. The results are shown in Fig. 8, which demonstrates that LASeR consistently selects the RM whose scoring aligns most closely with the ground-truth preferences.
@@
MAB Rewards Stabilize Over Time.
We observed empirically that the average return (i.e., normalized MAB rewards) of each iteration tends to stabilize over time, particularly after 6-8 training iterations. Fig. 9 shows per-RM normalized MAB reward over training iterations on GSM8K. Notably, the more effective RMs (e.g., Qwen and Zephyr for math reasoning tasks) retain higher MAB rewards over time, suggesting that convergence does not imply uniformity, but rather stabilization of relative effectiveness (this is also consistent with results in Fig. 6 and Fig. 7 that RM selection adapts to the best RM for each task).
@@
LASeR Benefits Cold-start Training.
As MAB rewards tend to converge over training iterations, the trained MAB on one dataset can be used for cold-start training on another dataset without retraining the MAB. We conducted an additional experiment where we reinitialized the LLM from the SFT checkpoint and re-trained on reasoning tasks using the converged MAB policy from a prior LASeR run on Wildchat. We denote this new method as cold-start LASeR (CS-LASeR) and present the result in the table below. Results show a comparable performance compared to LASeR (76.32% vs 76.24% on average of 3 tasks), suggesting that using a fixed, learned RM selector policy can benefit cold-start training.
@@
Detailed Results for Each RM.
Here, we provide detailed reasoning results for each chosen RM where we use a single RM during training (c.f. Sec. 3.1) in Table 6. These results demonstrate that Qwen1.5-7B-Chat outperforms other RMs on StrategyQA and MMLU, whereas on GSM8K Zephyr-7B-Alpha has the best performance with Llama-3-8B. However, LASeR still yields the best performance, outperforming all RMs by at least $1\%$ on average across reasoning tasks, without the knowledge of which RM is most suited for each task a priori. This is consistent with the results on Wildchat in Table 7 where LASeR outperforms all RMs with significant win rates (e.g., 63.47% and 59.02% against Qwen and Eurus, respectively).
@@
Impact of RMs Diversity.
In this experiment, we test the impact of RMs diversity on downstream performance. Instead of selecting RMs based on overall RewardBench scores across categories as in our initial experiments, we created a more capability-diverse RM set by choosing the best-performing RM in each category in RewardBench: "Chat", "Chat Hard", "Safety", and "Reasoning". This selection improves the diversity in the types of tasks and evaluation criteria each RM specializes in. The new RM set includes Qwen1.5-7B-Chat, Zephyr-7B-Alpha, Tulu-2-DPO-7B, and MMPO-Gemma-7B, replacing Eurus and OLMo models from the previous RM set. Using this updated RM set, we test LASeR’s ability (denoted as "LASeR-Diverse") to generalize across diverse categories on WildChat. We compare LASeR-Diverse to the Best RM (from the revised RM set) and RM Agreement Ensemble baselines (the strongest baselines across our benchmarks) and report the results in Table 8. These results show that with a new set of RMs, LASeR continues to perform better, outperforming both the Best RM and RM Agreement Ensemble baselines. We emphasize that whether RMs are selected based on overall scores or diversity across RewardBench categories, LASeR consistently outperforms baselines. This reflects the core contribution of LASeR: its ability to automatically select reward models at the instance level, achieving consistent gains across tasks without requiring manual tuning or oracle selection. Users can simply provide a pool of RMs without knowing which is best.

### v3

@@ Acknowledgements
We thank the reviewers and Justin Chih-Yao Chen for their useful feedback. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, the Accelerating Foundation Models Research Program, NSF CAREER Award 1846185, Apple PhD Fellowship, and Capital One Research Award. The views contained in this article are those of the authors and not of the funding agency.
@@ A.1 Experimental Setting
Training Setup.  We provide the training details for our method and baselines as follows:
@@
  • Batch size: We fine-tune the model using a learning rate of 5e65e\!-\!6 and a batch size of 16. Noted that we experimented with different batch sizes to evaluate their impact. Using a batch size of 1 yielded comparable performance to a batch size of 16 but was significantly less efficient in training the LLM due to the increased computational overhead. Based on this, we opted to use a batch size of 16 for a better trade-off between performance and efficiency. For datasets like Wildchat, which contain clearly defined and diverse categories, we structure the batches such that each batch consists of data belonging to a single category. This setup minimizes the risk of mismatches between the RM and the batch data, as each RM is evaluated on its most relevant data category. For reasoning datasets where such predefined categories do not exist, shuffling the data during training ensures diverse data and a good training signal for LLM within each batch. Even if individual batches vary, the repeated exposure to diverse inputs allows the bandit to learn which RMs are more helpful overall. Over training iterations, this appears to provide a robust signal for the bandit to learn effective RM selection, as demonstrated by LASeR’s consistent gains across datasets and the convergence behavior of MAB rewards shown in Appendix˜B.
@@
To test the hypothesis that the log probabilities of the chosen and rejected responses from the policy model are generally close in value, with the chosen response tending to have a slightly higher log probability, we conduct an experiment using Llama-3-8B and RewardBench dataset. For each instance, we calculated the reward difference by subtracting the log probability (reward) of the chosen response from that of the rejected response. Formally, given a policy model $\pi_{0}$, a prompt $x_{i}$ and a pair of chosen response $y_{i}^{w}$ and rejected response $y_{i}^{l}$ from RewardBench data, we calculate the reward difference as $\log\pi_{0}(y_{i}^{w}\mid x_{i})-\log\pi_{0}(y_{i}^{l}\mid x_{i})$. As shown in Fig.˜5, the distribution is centered near zero, indicating that the values are often close. However, the mass is skewed slightly to the right, suggesting that the chosen response tends to have a higher log probability than the rejected one in the majority of cases. This suggests the policy model generally has a reasonable ranking, but benefits from DPO’s maximization of the margin.
@@
Datasets.  For StrategyQA, GSM8K, and MMLU, we divided each dataset into training and test sets. The model is fine-tuned on the training set and dev set and then evaluated on the test set. For WildChat, the dataset was split into a 70/10/20 ratio for training, development, and testing. Following Zhao et al. [2024], prompt categorization is done using a small off-the-shelf classifier. [^5] Link: https://huggingface.co/valpy/prompt-classification For LongBench, we subsample 5K examples for three tasks: multi-document QA, summarization, and few-shot learning. Each category was split into a 70/10/20 ratio, and the bandit model was trained and validated on the training and development sets and then tested on the test set. We report the detailed number of instances for train, development, and test sets in Table˜3.
@@
Additional Results and Analysis.  In Appendix˜B, we present additional analyses on the statistical stability of LASeR and baselines across multiple runs (Table˜4). We show that the RM selected by LASeR adapts to input queries (Fig.˜6) and aligns well with RewardBench ground truth rankings (Fig.˜8). We also show that MAB rewards converge over training (Fig.˜9), and that the learned MAB policy can be reused to improve cold-start training (Table˜6). Detailed comparisons against individual RMs used in isolation (Tables˜8 and 9) further confirm that LASeR consistently outperforms all single-RM baselines. In addition, we show that LASeR maintains strong performance across different sets of RMs (Table˜10).
In Appendix˜C, we evaluate the generalization of LASeR across multiple dimensions: different bandit algorithms (Table˜11), base models and RMs (Tables˜12 and 14), number of RMs (Fig.˜11), training loss functions (Table˜17), and out-of-distribution tasks (Table˜18). We also show that LASeR can successfully select evaluation metrics tailored to domain-specific reasoning tasks [Golovneva et al., 2022, Prasad et al., 2023] (Table˜19). Finally, LASeR demonstrates robustness to both noisy RMs and underperforming reward signals (Figs.˜12 and 13).
@@ A.2 Details of Bandit Algorithms
Algorithm for Sec.˜3.3.  We provide the detailed algorithm for Sec.˜3.3 in Algorithm˜1.
@@
1:Input: LLM $\mathcal{M}$, reward models $\mathcal{R}=\{R_{1},R_{2},\dots,R_{K}\}$, dataset $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$, bandit algorithm (LinUCB)
@@
The score for arm $a_{t}$ is updated based on the observed normalized reward $\hat{\mathcal{L}}^{m}(t)$ and the probability $p_{k}(t)$ of selecting that arm:
@@
MAB reward normalization.  To maintain a consistent scale and magnitude of MAB rewards across training, we apply scaled rewards based on the quantiles of the reward history, following the method outlined by Graves et al. [2017]. Let $L=\{-\mathcal{L}^{m}(1),\dots,-\mathcal{L}^{m}(t-1)\}$ represent the unscaled reward history up to time step $t$. This history’s lower and upper quantiles are denoted as $q_{t}^{lo}$ and $q_{t}^{hi}$, respectively. We set $q_{t}^{\text{lo}}$ and $q_{t}^{\text{hi}}$ to be $20^{th}$ and $80^{th}$ quantiles. The scaled reward, $\hat{\mathcal{L}}^{m}(t)$, becomes:
[…]
@@ Appendix B Additional Empirical Results
Analysis on Statistical Significance of LASeR.  While our main results are based on single runs, we note that this is standard practice in iterative preference optimization setups, including recent works such as Pang et al. [2024], where large-scale LLM training limits the feasibility of multi-seed replication across all benchmarks. Nonetheless, to provide some insights regarding the statistical significance of LASeR, we report the mean and standard deviation over 5 seeds on reasoning tasks. We compare LASeR against the Best RM and RM Agreement Ensemble (identical random seeds to ensure fair comparison), which are the most competitive across our benchmarks. Table˜4 shows that LASeR consistently outperforms the baselines across runs, with standard deviations below 0.3% (lower than RM Agreement Ensemble), showing stable performance. We believe these consistent gains across tasks and models support the claim that LASeR is not only superior but also reliable in expectation. To further support these results, we conduct a non-parametric hypothesis test. Specifically, we evaluate the following hypotheses: Null hypothesis: LASeR’s accuracy is equal to that of the competing method; Alternative hypothesis: LASeR’s accuracy is greater than that of the competing method. We use a one-sided non-parametric Wilcoxon signed-rank test on paired accuracy values across random seeds. The p-values of the tests between LASeR and the two strongest baselines are reported in Table˜5. At a significance level of 0.05, the results in Table˜5 indicate that we can reject the null hypothesis, demonstrating that LASeR significantly outperforms both baselines across all three tasks.
@@
Table 5: Statistical significance (p-values) of LASeR and baselines on reasoning tasks.
LASeR vs. Baseline StrategyQA GSM8K MMLU
LASeR vs. Best RM 0.036 7e-4 0.019
LASeR vs. RM Agree. Ensemble 8e-3 0.024 0.046
LASeR’s Selected RM Adjusts to the Query.Fig.˜6 shows the relative utilization rates of each arm (i.e., RM) of the bandit on WildChat. We observed vastly different RM utilization rates depending on the underlying query within the same dataset. On queries requiring creativity in LLM responses, we find that Olmo and Eurus RMs are utilized about $20\%$ more often than Qwen RM, despite Qwen RM being ranked higher on RewardBench. This can be explained by the fact that the Qwen RM largely underperforms on the “chat” subsplit of RewardBench (behind Olmo and Eurus by nearly 40 points in chat score). On the other hand, Qwen RM is used roughly half the time for user prompts involving math, while Olmo and Eurus are used sparingly. This is consistent with Qwen RM’s ranking on the “reasoning” split of RewardBench, outperforming Eurus and Olmo RMs by 15-20 points. Note that LASeR automatically deduces these relative rankings of RMs and uses them depending on the underlying query without having access to the RewardBench leaderboard. Therefore, RM utilization of LASeR can serve as an analysis tool for future work when assessing performance on untested domains.
@@
To further validate our findings, we conducted an additional experiment on RewardBench data. In this setup, we used the MAB trained on Wildchat to select the most appropriate RM for each query. This selected RM was then used to choose the preferred response, which we compared against the ground-truth preferences in RewardBench. We also report the accuracy of each RM in selecting the correct preferred responses. The results are shown in Fig.˜8, which demonstrates that LASeR consistently selects the RM whose scoring aligns most closely with the ground-truth preferences.
@@
MAB Rewards Stabilize Over Time.  We observed empirically that the average return (i.e., normalized MAB rewards) of each iteration tends to stabilize over time, particularly after 6-8 training iterations. Fig.˜9 shows per-RM normalized MAB reward over training iterations on GSM8K. Notably, the more effective RMs (e.g., Qwen and Zephyr for math reasoning tasks) retain higher MAB rewards over time, suggesting that convergence does not imply uniformity, but rather stabilization of relative effectiveness (this is also consistent with results in Fig.˜6 and Fig.˜7 that RM selection adapts to the best RM for each task).
@@
LASeR Benefits Cold-start Training and RM Adaptation.  As MAB rewards tend to converge over training iterations, the trained MAB on one dataset can be used for cold-start training on another dataset without retraining the MAB. We conducted an additional experiment where we reinitialized the LLM from the SFT checkpoint and re-trained on reasoning tasks using the converged MAB policy from a prior LASeR run on Wildchat. We denote this new method as cold-start LASeR (CS-LASeR) and present the result in the table below. Results show a comparable performance compared to LASeR (76.32% vs 76.24% on average of 3 tasks), suggesting that using a fixed, learned RM selector policy can benefit cold-start training.
@@
Another benefit of LASeR is that it can add or update an RM without requiring the entire MAB to be retrained from scratch. LinUCB maintains independent statistics $(A_{k},b_{k})$ for each RM arm. To add a new RM, we can initialize its parameters and begin exploration, allowing the bandit to learn its utility online. When adding a new RM to LASeR, only the new RM’s bandit parameters need to be initialized (e.g., $(A_{k}=I,b_{k}=0)$). The existing RMs’ statistics remain unchanged, meaning the MAB retains all prior learning about their utility. This allows the bandit to continue exploiting well-performing existing RMs while gradually exploring the new RM. Similarly, if an existing RM is updated (e.g., improved or retrained), the MAB can re-learn its statistics gradually through continued training. To empirically demonstrate this, we conduct an experiment where we resume training Llama-3-8B from a saved checkpoint and add a new, stronger RM QRM-Llama3-8B based on the latest RewardBench [Lambert et al., 2024] leaderboard to the current pool of RMs. We observe that the MAB quickly adapts to the new RM, with training converging after just 2 additional iterations and achieving 75.58% accuracy on the GSM8K test set, improving over LASeR without QRM-Llama3-8B by 0.83%. We also run an additional experiment where we reset the MAB parameters for an existing RM (Qwen) and then resume training Llama-3-8B from a saved checkpoint. Training converges after 3 additional iterations and achieves comparable accuracy to the original checkpoint before the reset (74.72% vs. 74.75% accuracy).
Besides accuracy, we also show the utilization rates of each RM before (Checkpoint), after (Transfer) adding the new RM, and restart (Restart) in Table˜7. These results show that LASeR can effectively integrate new RMs and adapt its selection strategy, shifting toward using stronger RMs when they offer more informative supervision, while still retaining and utilizing previously effective RMs like Qwen for reasoning tasks. Additionally, in our restart experiment, where we reset the MAB parameters of Qwen and resume training, demonstrates that LASeR can quickly relearn the utility of a strong RM, returning to similar usage levels with just a few additional iterations.
Table 7: Utilization rate (%) across RMs when including a different RM or restarting a RM.
Utilization Rate (%) Olmo Eurus Qwen Zephyr QRM
Checkpoint 6.26 2.06 66.81 24.87 -
Transfer (Add QRM) 1.68 0.77 35.08 8.93 53.54
Restart (Qwen) 7.68 2.47 65.25 24.59 -
Detailed Results for Each RM.  Here, we provide detailed reasoning results for each chosen RM where we use a single RM during training (c.f. Sec.˜3.1) in Table˜8. These results demonstrate that Qwen1.5-7B-Chat outperforms other RMs on StrategyQA and MMLU, whereas on GSM8K Zephyr-7B-Alpha has the best performance with Llama-3-8B. However, LASeR still yields the best performance, outperforming all RMs by at least $1\%$ on average across reasoning tasks, without the knowledge of which RM is most suited for each task a priori. This is consistent with the results on Wildchat in Table˜9 where LASeR outperforms all RMs with significant win rates (e.g., 63.47% and 59.02% against Qwen and Eurus, respectively).
@@
Impact of RMs Diversity.  In this experiment, we test the impact of RMs diversity on downstream performance. Instead of selecting RMs based on overall RewardBench scores across categories as in our initial experiments, we created a more capability-diverse RM set by choosing the best-performing RM in each category in RewardBench: "Chat", "Chat Hard", "Safety", and "Reasoning". This selection improves the diversity in the types of tasks and evaluation criteria each RM specializes in. The new RM set includes Qwen1.5-7B-Chat, Zephyr-7B-Alpha, Tulu-2-DPO-7B, and MMPO-Gemma-7B, replacing Eurus and OLMo models from the previous RM set. Using this updated RM set, we test LASeR’s ability (denoted as "LASeR-Diverse") to generalize across diverse categories on WildChat. We compare LASeR-Diverse to the Best RM (from the revised RM set) and RM Agreement Ensemble baselines (the strongest baselines across our benchmarks) and report the results in Table˜10. These results show that with a new set of RMs, LASeR continues to perform better, outperforming both the Best RM and RM Agreement Ensemble baselines. We emphasize that whether RMs are selected based on overall scores or diversity across RewardBench categories, LASeR consistently outperforms baselines. This reflects the core contribution of LASeR: its ability to automatically select reward models at the instance level, achieving consistent gains across tasks without requiring manual tuning or oracle selection. Users can simply provide a pool of RMs without knowing which is best.